OpenStudy (anonymous):

The length of a rectangle is 3 meters more than its width. If the area is 108 square meters, what are the length and width?

OpenStudy (anonymous):

x = y+3 x = length y = width (y+3)(y) = 108 y^2 +3y -108 = 0 y = 9 x = 12

OpenStudy (anonymous):

let breadth be x m then length is x+3m x(x+3)=108 x^2+3x=108 solving this equation we get breadth(x)=9 and length(x+3)=12